OpenStudy (vortish):

evaluate the function f(x)=X^2+8x+9 at the given values of the independent variable and simplfi a) f(3) B)F(x+2) c)f(-x)

OpenStudy (jhannybean):

So you've got your function \(\ f(x)=x^2+8x+9\) and your problem is basically telling you when you have values of \(\ x= 3, x+2, -x\) what would your function be? All we have to is plug in all these x-values in place of your function and simplify it :)

OpenStudy (jhannybean):

let's start with (c) , when f(-x) = x^2 +8x+9 For this one, all we have to do is replace the x with -x and simplify it. f(-x) = (-x)^2 +8(-x) + 9 Can you simplify this for me?

OpenStudy (vortish):

x =3^2+8(3)+9 a=42

OpenStudy (jhannybean):

Yess good job. you got this! :)

OpenStudy (vortish):

now I am working on the b and it the equation is f(x+2)=(x+2)^2+8(x+2)+9

OpenStudy (jhannybean):

and remember, f(x) = y so essentially, f(x) is a function of x, therefore we can rewrite this problem as : y(x) = x^2 +8x +9

OpenStudy (jhannybean):

And you are absolutely correct for (b)

OpenStudy (jhannybean):

\(\ (x + y)^2 = x^2 +2xy + y^2\) , use this when expanding the first part o your function :D

OpenStudy (jhannybean):

I mean expand it using this form* if you were stuck on that!

OpenStudy (vortish):

figging hate this darn stuff cuz it just confuses me

OpenStudy (jhannybean):

What part is confusing you? I'll see if I can help you clarify that :D

OpenStudy (vortish):

the x+y were does y show up in the equation

OpenStudy (jhannybean):

Oh I meant \[\ (\color{blue}x + \color{green}y)^2 = (\color{blue}x + \color{green}2)^2\] I was showing you that if you followed my expansion for (x+y)^2 you can use the same method for expanding (x+2)^2 :)

OpenStudy (jhannybean):

and i'm just setting these two equal to eachother to show how the equations math up, not saying they are equal to eachother., haha.

OpenStudy (vortish):

ok I understand now

OpenStudy (jhannybean):

Maybe I should have written it... \[\ (\color{blue}x + \color{green}y)^2 =\]\[\ (\color{blue}x + \color{green}2)^2=\] Yeah that's what I meant when I was saying a comparison, haha,.

OpenStudy (vortish):

ok ty sorry I get really frustrated with complex forumlas

OpenStudy (jhannybean):

Oh no problem. What did you get for the (b) and (c)?

OpenStudy (vortish):

im still tryi8ng to figure out b

OpenStudy (jhannybean):

Ahh, what do you have so far for b? :3

OpenStudy (vortish):

2that the (x+2)^2 is (x+2)(x+2)+8(x+2)+9

OpenStudy (jhannybean):

ok, and if (x+2)^2 = (x+2)(x+2), do you know how to use the foiling method? :o

OpenStudy (vortish):

i forgotten it long ago and had trouble with it when i used it in my last math class a year ago

OpenStudy (jhannybean):

Ah, ok. FOIL stands for First, Outter, Inner, Last. Let's approach this step by step.

OpenStudy (vortish):

so one would be 5x and the other is also 5x right

OpenStudy (jhannybean):

Can you explain to me how you got the 5x?

OpenStudy (vortish):

not five I am looking at the example it would be 2x and 2x sorry

OpenStudy (jhannybean):

Well, I guess that is part of it. :) you're on the right track! Let's approach this using the FOIL method so if you come across another polynomial of this type, you'll understand how to properly expand it :D

OpenStudy (jhannybean):

\(\ (\color{blue}x + \color{green}2 )^2 = (\color{blue}x + \color{green}2)(\color{blue}x + \color{green}2)\) 1. First : \(\ (\color{blue}x +2)(\color{blue}x +2) = \color{red}{x^2}\) 2. Outter : \(\ (\color{blue}x +2)( x + \color{green}2)= \color{red}{2x}\) 3. Inner : \(\ (x + \color{green}2)(\color{blue}x + 2) = \color{red}{2x}\) 4. Last: \(\ (x +\color{green}2)(x+\color{green}2) = \color{red}4\) Adding these all up together, we get: \(\ \color{red}{x^2 + (2x + 2x) + 4} = x^2 + 4x + 4\)

OpenStudy (jhannybean):

Now following the colors and the acronym FOIL, can you see what multiplies with what and how I get my results? :o

OpenStudy (vortish):

but its not 4 its 8

OpenStudy (jhannybean):

Which part are you referring to? Sorry, i'm a bit confused D:

OpenStudy (vortish):

you said it s x^2+(2x+2x)+4 when the 4 should of been 8x I think

OpenStudy (jhannybean):

Nope! :D I just helped you expand the first PART of your function, when f(x+2) = x^2 So when you plug in (x+2) for x^2 you get (x+2)^2 and (x+2)^2 = x^2 +4x + 4 :D

OpenStudy (jhannybean):

Now you've for your other 2 parts, the 8(x+2) + 9 :P

OpenStudy (vortish):

would the equation look like 2x+2x+8x+16+9

OpenStudy (jhannybean):

I don't think so D:

OpenStudy (jhannybean):

Well, let's see. do you understand how (x+2)^2 = x^2 +4x + 4 came to be when we expanded the polynomial? If you follow the foil method I posted a few posts up, that should help a bit!

OpenStudy (vortish):

not really I see how you do it but dont understand where you get the four terms from

OpenStudy (jhannybean):

Hmm..alright. The first term comes when we multiply the "first" terms of EACH multiplier. that is, we multiply the x to the x. Are you with me? :D

OpenStudy (vortish):

yes

OpenStudy (jhannybean):

Ok, that is our FIRST term. Our second term comes when multiply the "outer" numbers. We take the "x" from our first factor, and multiply it to the "outer" number, "2", therefore we get 2x. Let me know when you understand this as well :3

OpenStudy (vortish):

i understand how you get the 2x+2x but from there is were i get lost

OpenStudy (jhannybean):

Oh, add the 2x + 2x together because they are like-terms :D therefore you get 4x!

OpenStudy (jhannybean):

I put parenthesis () around them to group them together, letting you know they were like-terms and needed to be added together :P

OpenStudy (vortish):

ok so that is were we get the x^2+4x+4 from right so were does the 8(x+2)+9 come in

OpenStudy (vortish):

i am really sorry that you are having to give up so mcuh time to help me with this equation

OpenStudy (jhannybean):

because your main function is f(x) = x^2 + 8x+ 9, we only solved the x^2 part. You're evaluating your main function when f(x) = f(x+2), so now you have to expand 8(x+2)

OpenStudy (jhannybean):

Oh no! it's ok, don't apologize xD as long as you understand how to solve other questions like this I'll be happy :)

OpenStudy (jhannybean):

Basically we replaced ALL your ORIGINAL x's with (x+2)'s.

OpenStudy (vortish):

the 8(x+2) would be 8x+16 right

OpenStudy (jhannybean):

Mmhmm :)

OpenStudy (vortish):

so the whole equation would look like x^2+4x+4+8x+16+9 and combine like terms

OpenStudy (jhannybean):

Yes! you got it :D

OpenStudy (vortish):

so the answer is X^2+12x+29

OpenStudy (jhannybean):

You got it :)

OpenStudy (vortish):

do you have time to help me with part c

OpenStudy (jhannybean):

For (c) we're just replacing all the x's in your original function with -x

OpenStudy (jhannybean):

so it'll be f(-x) = (-x)^2 + 8(-x) + 9 simplify this and you'll have your answer here too :)

OpenStudy (vortish):

ok so (-x)(-x)_8x+9 correct

OpenStudy (vortish):

(-x)(-x)-8x+9

OpenStudy (jhannybean):

mmhmm, and when you multiply (-x)(-x) you get?

OpenStudy (vortish):

+x

OpenStudy (vortish):

x^2-8x+9

OpenStudy (jhannybean):

You got it :)

OpenStudy (vortish):

ty I tryed to fan you but I can not right now I will when I can

OpenStudy (jhannybean):

Thanks! :D